Reversible Computing Bibliography

MIT Reversible Computing Project Artificial Intelligence Lab 545 Technology Sq. Cambridge, Massachusetts 02139

Typeset on January 21, 1998

References

[1] William C. Athas, Lars “J.” Svensson, Jeffrey G. Koller, Nestoras Tzartzanis, and Eric Ying-Chin Chou. Low-power digital systems based on adiabatic-switching principles. IEEE Transactions on Very Large Scale Integration (VLSI) Systems, 2(4):398–407, December 1994.

[2] Stephen Avery and Marwan Jabri. Design of a register file using adiabatic logic. Technical Report SCA-06/06/97, University of Sydney SEDAL, June 1997.

[3] Henry G. Baker. Lively linear lisp—‘Look ma, no garbage!’. ACM SigPlan Notices, 27(8):89–98, August 1992.

[4] Henry G. Baker. NREVERSAL of fortune — the thermodynamics of garbage collection. In Y. Bekkers, editor, International Workshop on Memory Management, pages 507–524. Springer-Verlag, 1992.

[5] Theodore Baker, John Gill, and Robert Solovay. Relativizations of the P =?N P question. SIAM J. Computing, 4(4):431–442, December 1975.

[6] Edward Barton. A reversible computer using conservative logic. Term paper for 6.895 at MIT, 1978.

[7] Jacob D. Bekenstein. Universal upper bound on entropy-to-energy ratio for bounded systems. Phys. Rev. D, 23(2):287–298, January 1981.

[8] Jacob D. Bekenstein. Entropy content and information flow in systems with limited energy. Phyical Review D, 30(8):1669–1679, 15 October 1984.

[9] C. H. Bennett. Logical reversibility of computation. IBM J. Research and Development, 17(6):525–532, 1973.

[10] C. H. Bennett. The thermodynamics of computation, a review. Int’l J. Theoretical Physics, 21(12):905–940, 1982.